OpenStudy (anonymous):

Simplify the expression. State any excluded values. Show your work. 2a^2-4a+2 / 3a^2-3

OpenStudy (anonymous):

Its 2(a−1)/3(a+1) when simplified but i need to write it step by step. And the last question

OpenStudy (anonymous):

Thanks

OpenStudy (solomonzelman):

2a^2-4a+2 = 2(a^2-2a+1) = 2(a-1)^2 3a^2-3 = 3(a^2-1) = 3(a-1)(a+1) So, let's rewrite the question 2(a-1)^2 / 3(a-1)(a+1) cancel out (a-1) 2(a-1) / 3(a+1)

OpenStudy (anonymous):

Thats all it want me to do then? and thanks for taking the time out to show me step by step

OpenStudy (solomonzelman):

np it wasn't hard for me ;)

OpenStudy (anonymous):

Wow, must be smart! i hate math i have 4 more problems like this but difficult for me.

OpenStudy (anonymous):

@SolomonZelman see i also gota do this on step by step. this other one says, multiply or divide, x^2+13x+40/ x-7 (Divided by) x+8/x^2-49